Natures's Best wants to combine nuts they sell for $3.60 a pound with dried fruit they sell for 2.40 a pound to create a trail mix. How much of each snack should they use to make 10 pounds of trail mix that would sell for $3.30 por pound?

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

x*3.60+(10-x)*2.40=10*3.3

Answer by stanbon(75887)

Natures's Best wants to combine nuts they sell for $3.60 a pound with dried fruit they sell for 2.40 a pound to create a trail mix.
--------------------------------------------------------------
How much of each snack should they use to make 10 pounds of trail mix that would sell for $3.30 por pound?
---
Equations:
Quantity Eq.::: n + f = 10
Value Eq.::: 360n+ 240f = 330*10
-----------------------------------
Multiply thru the Quantity Equation by 360 and solve for "f":
360n + 360f = 3600
360n + 240f = 3300
------------------------
120f = 300
f = 2.5 lbs (# of lbs. of fruit in the mixture)
----
Since n+f = 10, n = 7.5 lb. (# of lbs. of nuts in the mixture)
========================================
Cheers,
Stan H.